OpenStudy (anonymous):

Correct the following samples of gases to standard temp. ( assume pressure is constant). STP stands for standard temp and pressure which is 273K &101.3kPa. T2 is 273K and on last three P2 is 101.3kPa. 1) 1622cm3 at 15.0 C 2) 12.33L at 280K 3) 155.6 dm3 at 110K4 4) 1250cm3 at-45.0C 5) A sample of neon is collected at 1273.51 kPa and 12.0C. It has a volume of 2.25 L. What would the volume be of this gas at STP? 6) A sample of Argon has a volume of 1205 cm3 when its temp is -44.0C& its pressure is .937kPa. What would be the volume of the argon at STP?

OpenStudy (aaronq):

you're using the same equation \[\frac{ V _{1} }{ T _{1} }=\frac{ V _{2} }{ T _{2} }\] when volume and temp. or when pressure and temp, just substitute P instead of V in the same place. it works because PV=nRT -> P/T=nR/V if n,R.V are constant

OpenStudy (anonymous):

would it look like this1622/15.0?

OpenStudy (aaronq):

it needs to be in kelvin !

OpenStudy (aaronq):

1622cm^3/(15.0+273)K=V/273K V=1622cm^3*273K/(15.0+273)K

OpenStudy (anonymous):

hey I think I have this finally;D can check my answers 1) 1536.72cm^3 2) 12.646L 3) 386.17dm^3 4)1496cm^3 5)27.1L 6)13.29cm^3